OpenStudy (anonymous):

Calculate the PH of 0.175 M HY, if Ka=1.50x10^(-4)

OpenStudy (aaronq):

so write an ionization expression: \(Ka=\dfrac{[H^+][X^-]}{[HX]}\) make an ICE table, plug Ka and the values from E into the expression, solve for x. in this example x=\([H^+]\), use \(pH=-log[H^+]\)

OpenStudy (anonymous):

hi @aaronq, thank you. I realized i need to use the quadratic formula but for some reason my x values aren't giving me the right answer :|

OpenStudy (aaronq):

did you get \([H^+]\)=0.005 M?

OpenStudy (anonymous):

@aaronq somehow i got something totally different :/

OpenStudy (anonymous):

how did you get your answer?

OpenStudy (aaronq):

was this what you had? \(1.50*10^{-4}=\dfrac{x^2}{(0.175 -x)}\)

OpenStudy (anonymous):

yes ! @aaronq

OpenStudy (aaronq):

hm it's likely you made mistakes solving it, try again? if you dont get the right values i'll show you in steps.

OpenStudy (anonymous):

hm :/ I used the quadratic formula, is that what you had used as well?

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

yeah, the set up was: \(x=\dfrac{\sqrt{42009}-3}{40000}\)

OpenStudy (anonymous):

i had it set up this way...:\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] \[\frac{ -(1.50\times10^{-4}) \pm \sqrt{(1.50\times10^{-4}) ^{2}-4(1)(2.626\times10^{-5})} }{ 2 }\]

OpenStudy (aaronq):

i did it with wolfram, hold on i'll double check it

OpenStudy (anonymous):

ok! this is where i was most stuck :|

OpenStudy (aaronq):

ohhh you missed a negative sign, it's \(-2.626*10^{-5}\)

OpenStudy (anonymous):

did i!? thats weird b/c i did try it with the negative as well and it didn't work for some reason. I think it may just be my work :/

OpenStudy (aaronq):

should work !

OpenStudy (aaronq):

it's easy to screw up the quadratic equation, dont feel bad lol

OpenStudy (anonymous):

Ok @aaronq ill give it another try but i have the right approach correct?

OpenStudy (aaronq):

Yeah, the exact equation is \(x=\dfrac{(-1.50x10^{-4})+\sqrt{(1.50x10^{-4})^2-4*(-2.625*10^{-5}})}{2}\)

OpenStudy (anonymous):

thank you so much! @aaronq

OpenStudy (aaronq):

no problem, glad you got that sorted out!